24. Draw the polymer formed from the monomer vinyl chloride, CH,CHCI:
Answer #1

CH2=CH-Cl is vinyl chloride

some many vinyl chloride molecules are under goes polymerization to polyvinylchloride (PVC) at high pressure and temperature.

CH2=CH-Cl is monomer   

n(CH2=CH-Cl) ---------------------> -[CH2-CH-Cl-]n

-[CH2-CH-Cl-]n is polymer